# A Survey on Privacy-Preserving Caching at Network Edge: Classification, Solutions, and Challenges

### v1

@@ Abstract.
Caching content at the network edge is a popular and effective technique widely deployed to alleviate the burden of network backhaul, shorten service delay and improve service quality. However, there has been some controversy over privacy violations in caching content at the network edge. On the one hand, the multi-access open edge network provides an ideal surface for external attackers to obtain private data from the edge cache by extracting sensitive information. On the other hand, privacy can be infringed by curious edge caching providers through caching trace analysis targeting to achieve better caching performance or higher profits. Therefore, an in-depth understanding of privacy issues in edge caching networks is vital and indispensable for creating a privacy-preserving caching service at the network edge. In this article, we are among the first to fill in this gap by examining privacy-preserving techniques for caching content at the network edge. Firstly, we provide an introduction to the background of Privacy-Preserving Edge Caching (PPEC). Next, we summarize the key privacy issues and present a taxonomy for caching at the network edge from the perspective of private data. Additionally, we conduct a retrospective review of the state-of-the-art countermeasures against privacy leakage from content caching at the network edge. Finally, we conclude the survey and envision challenges for future research.
@@ 1. Introduction
Content caching at the network edge is driven by two factors. First, the population of networked devices has become astronomical due to advances in intelligent terminals and the broad deployment of the Internet of Things (IoT) (Cui et al., 2022; Guo et al., 2022; Cui et al., 2023). Second, the Internet content market is blooming due to the proliferation of various multimedia content (Zhang et al., 2022d; Ni et al., 2021). As a result, network-based content delivery services are extremely bandwidth-consuming. It was forecasted by Cisco (others Forecast, 2019) that the consumer share of the total devices, including both fixed and mobile devices, will be more than 21 billion and account for 74% of total devices in 2023. At the same time, emerging network technologies, such as Gigabit Ethernet and 5G, are expected to provide extremely high data transmission rates and low access delays for terminal devices at the network edge to support time-sensitive services such as autonomous driving, industrial automation, high-quality video streaming, and virtual/enhanced emerging applications.
@@
Such a vast data flow brings two main challenges to the established networks: (1) It brings a heavy communication burden to the Internet core network links. During the peak hours of network usage, a large amount of content transmission will inevitably aggravate the link burden of the core network, causing network congestion and the increase of network operating costs; (2) It will also prolong the service delay of content transmission from remote servers to end devices, which will adversely influence users’ service Quality-of-Experience (QoE) or even ruin the reliability of delay-sensitive applications.
@@
Edge Caching is a technique that involves storing content in proximity to end users, typically at or near the point of user access or ahead of the core network (Ni et al., 2021). Its primary objective is to shorten service latency and enhance content delivery performance by bringing content closer to the users who request it. When users request content that is available in the edge cache, their requests can be directly served at the network edge with a high Quality-of-Service (QoS). However, if the requested content is not available in the edge cache, it can be redirected to a remote server, such as a data center. In a typical edge network (Ni et al., 2021), there are three main entities for edge caching: end devices, access networks and edge networks:
  1. (1)
    End devices (e.g., smartphones, laptops, intelligent vehicles, and industrial IoT devices) carried by users will generate requests for downloading content via networks (Cui et al., 2022; Guo et al., 2022). It is possible that end devices can share content with each other through Device-to-Device (D2D) communications with licensed-band or unlicensed-band protocols.
  2. (2)
    In access networks, wired and wireless communication technologies support end devices for accessing the Internet with the infrastructures in the edge network, such as 5G Base Stations (BSs), WiFi routers, intelligent television boxes, and Roadside Units (RSU) in the Internet of Vehicles (IoV). Popular content can be cached in these accessible infrastructures at the instance of end devices or third-party Internet Service Providers (ISPs) to serve user requests.
  3. (3)
    Edge cache can be located at Edge Servers (ESs) ahead of the core network, such as nodes in the Content Delivery Network (CDN) (Cui et al., 2020a), edge routers in the Information-Centric Network (ICN) (Sivaraman and Sikdar, 2021; Xue et al., 2019, 2018), and macro base stations covering a specific region (Araldo et al., 2018). We also call it in-network-edge caching. Edge servers, generally maintained by third-party suppliers, are the heart and anchor point for multi-access edge networks to enhance various content delivery applications.
@@
Caching content at the edge network is effective in reducing the burden of network backhaul (Yang et al., 2019; Jiang et al., 2017; Qiao et al., 2022), shortening service latency (Zhang et al., 2022b; Qiao et al., 2022; Cui et al., 2020a), and diminishing resource cost (Jiang et al., 2017; Hassanpour et al., 2023). First, it is common to cache popular content at the edge network through which the edge network can offload the access of requests and hence reduce the backhaul data flow. Even though the caching capability is limited at the network edge, the edge cache can offload up to 35% of the traffic burden over backhaul links  (Ni et al., 2021). Second, the service latency can be significantly shortened by caching content on edge devices near end users. In particular, a shortened latency is critical for content delivery of latency-sensitive applications (Ni et al., 2021). Third, network edge can make the access of content inexpensive since caching content at edge devices can avoid the transmission bottleneck. For example, in wireless edge networks, spectral efficiency and energy efficiency can be improved by about 900% and 500%, respectively, by making use of network edge for caching content (Liu et al., 2016).
@@
Despite the enormous benefits brought by caching content at the network edge, there have been some controversy over privacy violations brought by such caching. The concerns can be illustrated from two aspects. The first privacy threat comes from external attackers, such as malicious user devices (Sivaraman and Sikdar, 2021; Liang and Liu, 2019; Acs et al., 2019; Qian et al., 2020; Cui et al., 2020a; Tong et al., 2022). The multi-access open property of the edge network provides an ideal surface for external attackers to obtain the cached content from the edge cache so as to extract sensitive information of end users (Xu et al., 2020). In particular, adversaries can obtain user-sensitive information by launching cache side-channel attacks (Sivaraman and Sikdar, 2021; Liang and Liu, 2019; Acs et al., 2019) and cache tampering attacks (Qian et al., 2020; Cui et al., 2020a; Tong et al., 2022). However, it is non-trivial to embed advanced privacy protection mechanisms into network edge due to the limited computing capacity, energy power, and storage space of edge devices. Second, user privacy can be infringed by curious edge caching providers by analyzing traces and management records. Due to limited caching space relative to the rapidly growing user population and the scale of content (Zhou et al., 2019), network edge providers have a strong motivation to spy on user privacy in order to improve their resource utilization. In other words, if content popularity can be accurately predicted, the right content can be cached by edge devices just before the surge of requests towards these content (Zhang et al., 2022d). Hence, edge network providers are curious about users’ personal interests and confidential information to infer their request behaviours, which can be extracted from users’ historical request traces (e.g., request patterns (Zhang et al., 2022d; Cui et al., 2020c, a), identifiable information (Zhang et al., 2022c; Zhu et al., 2021; Araldo et al., 2018; Cui et al., 2020c, a)). Network edge providers can implement monitoring attacks (Xue et al., 2018; Zhang et al., 2022c) and inference attacks (Qiao et al., 2022; Liu et al., 2022) in their systems to compromise users’ privacy based on collected request information from users. Therefore, an in-depth understanding of privacy risks in Privacy-Preserving Edge Caching (PPEC) is crucial for the design of feasible solutions to achieve privacy-preserving content cache at the network edge.
Recently, significant progress has been made by existing works to enhance the privacy protection for users in network edge (Ni et al., 2021), but it is still far from thoroughly solving all problems. Particularly, the influence of privacy protection on caching performance has been largely overlooked in existing works. In view of the limitations mentioned above and the scarcity of literature review on the edge-assisted network cache, this article is dedicated to comprehensively examining and categorizing current works on privacy issues at the edge caching networks. The main contributions of this article are summarized as follows:
  1. (1)
    We make in-depth discussions on sensitive privacy in edge caching and propose a taxonomy from a data perspective to classify existing works. To the best of our knowledge, this is the first such comprehensive exposition.
  2. (2)
    We conduct a thorough review of recent high-quality research, diving into the background of privacy attacks and mitigation methods in the realm of edge caching. Our review encompasses the latest solutions proposed for enhancing privacy in the edge cache, which have been published in influential conferences and journals in the fields of computing networks, architecture, and privacy, such as CCS, INFOCOM, ToN, JSAC, TPDS, TIFS, and TDSC, among others. Based on different kinds of privacy information and attacks towards each kind of privacy information, we respectively review countermeasures to defend against attacks for protecting each kind of infringed privacy.
  3. (3)
    Based on open problems outlined in existing works, we envision privacy-related open challenges in PPEC to provide insights for inspiring future research.

### v3

@@ Abstract.
Caching content at the edge network is a popular and effective technique widely deployed to alleviate the burden of network backhaul, shorten service delay and improve service quality. However, there has been some controversy over privacy violations in caching content at the edge network. On the one hand, the multi-access open edge network provides an ideal entrance or interface for external attackers to obtain private data from edge caches by extracting sensitive information. On the other hand, privacy can be infringed on by curious edge caching providers through caching trace analysis targeting the achievement of better caching performance or higher profits. Therefore, an in-depth understanding of privacy issues in edge caching networks is vital and indispensable for creating a privacy-preserving caching service at the edge network. In this article, we are among the first to fill this gap by examining privacy-preserving techniques for caching content at the edge network. Firstly, we provide an introduction to the background of privacy-preserving edge caching (PPEC). Next, we summarize the key privacy issues and present a taxonomy for caching at the edge network from the perspective of private information. Additionally, we conduct a retrospective review of the state-of-the-art countermeasures against privacy leakage from content caching at the edge network. Finally, we conclude the survey and envision challenges for future research.
@@ 1. Introduction
Content caching at the edge network is driven by two factors. First, the population of networked devices has become astronomical due to advances in intelligent terminals and the broad deployment of the Internet of Things (IoT) (Cui et al., 2022; Guo et al., 2022; Cui et al., 2023; Zhao et al., 2023). Second, the Internet content market is blooming due to the proliferation of various multimedia content (Zhang et al., 2022d; Ni et al., 2021). According to the report by Splunk, a Cisco company, there will be approximately 5.44 billion Internet users worldwide in 2024, including 5.07 billion social media users (Splunk, Inc., 2023). As a result, network-based content delivery services are extremely bandwidth-consuming. At the same time, emerging network technologies, such as Gigabit Ethernet, and 5G and beyond, are expected to provide extremely high data transmission rates and low access delays for terminal devices at the edge network to support time-sensitive services such as autonomous driving, industrial automation, high-quality video streaming, and virtual/enhanced emerging applications.
@@
Such a vast data flow brings two main challenges to the established networks: (1) It brings a heavy communication burden to the Internet core network links. During the peak hours of network usage, a large amount of content transmission will inevitably aggravate the link burden of the core network, causing network congestion and increasing network operating costs; (2) It will also prolong the service delay of content transmission from remote servers to end devices, which will adversely influence users’ service Quality-of-Experience (QoE) or even ruin the reliability of delay-sensitive applications.
@@
Edge Caching is a technique that involves storing content in close proximity to end users, typically at or near the point of user access or ahead of the core network (Ni et al., 2021). Its primary objective is to shorten service latency and enhance content delivery performance by bringing content closer to the users who request it. When users request content that is available in edge caches (ECs), their requests can be directly served at the edge network with a high Quality-of-Service (QoS). However, if the requested content is not available in the EC, it can be redirected to a remote server, such as a data center. Here, we make a brief introduction to edge caching from five aspects:
@@
Benefit of edge caching. Caching content at the edge network is effective in reducing the burden of network backhaul (Yang et al., 2019; Jiang et al., 2017; Qiao et al., 2022), shortening service latency (Zhang et al., 2022b; Qiao et al., 2022; Cui et al., 2020a), and diminishing resource cost (Jiang et al., 2017; Hassanpour et al., 2023). First, it is common to cache popular content at the edge network through which the edge network can offload the access of requests and hence reduce the backhaul data flow. Even though the caching capability is limited at the edge network, edge caches can offload up to 35% of the traffic burden over backhaul links  (Ni et al., 2021). Second, the service latency can be shortened by caching content on edge devices near end users. In particular, a shortened latency is critical for content delivery of latency-sensitive applications (Ni et al., 2021). Third, edge networks can make content access inexpensive since caching content at edge devices can avoid the bottleneck. For example, in wireless edge networks, spectral efficiency and energy efficiency can be improved by about 900% and 500%, respectively, by using edge networks for caching content (Liu et al., 2016).
Where to cache: Building on the work of Ni et al. (Ni et al., 2021), we further identify three main entities in edge networks for edge caching as follows: (1) End devices (e.g., smartphones, laptops, intelligent vehicles, and industrial IoT devices) carried by users will generate requests for downloading content via networks (Cui et al., 2022; Guo et al., 2022). It is possible that end devices can share content through Device-to-Device (D2D) communications with licensed-band or unlicensed-band protocols. (2) Access infrastructures, utilizing wired and/or wireless communication technologies, can support end devices in accessing the Internet. These infrastructures include 5G small base stations (SBSs) (Xu et al., 2019), WiFi routers, local switches, and roadside units (RSUs) in the Internet of Vehicles (IoV) (Cui et al., 2020c; Zhang et al., 2022c). Popular content can be cached within these access infrastructures to promptly serve user requests. (3) Edge servers (ESs) positioned ahead of the core network, such as edge nodes (ENs) in the Content Delivery Network (CDN) (Cui et al., 2020a), edge routers in the Information-Centric Network (ICN) (Sivaraman and Sikdar, 2021; Xue et al., 2019, 2018), and macro base stations operated by Internet service providers (ISPs) (Araldo et al., 2018), can be utilized as ECs, a concept known as in-network-edge caching. These ESs, typically maintained by third-party suppliers, are the core points for multi-access edge networks, enhancing various content delivery applications.
What to Cache. In edge caching systems, determining what content to cache is crucial for optimizing cache space utilization and reducing latency. The content to be cached generally falls into three categories: (1) User-related popular content: This includes content that is frequently requested by end users, such as web pages, videos, images, and other multimedia files (Cui et al., 2020a; Wu et al., 2016). Caching such content at the edge improves user experience by reducing service delay when users access commonly accessed content. (2) Public and static content: This category includes high-reuse, non-user-specific content associated with applications (Ni et al., 2021), such as JavaScript files, CSS stylesheets, icons, PDF documents, and API responses. Caching these static resources decreases application load times and reduces the burden on central servers. (3) Edge-computable and storable content: This includes data that can be computed and stored directly at the edge, such as model parameters for federated learning (Liu et al., 2022; Qiao et al., 2022), user patterns and content popularity for edge caching decisions (Cui et al., 2020c, a), and IoT sensor data awaiting processing (Wang et al., 2020; Yu et al., 2022). Caching such content helps minimize backhaul traffic, enables efficient edge processing, and reduces overall latency by avoiding redundant computations.
How to Cache. Edge caching strategies can be broadly classified into reactive and proactive approaches. (1) Reactive caching employs eviction-based methods that decide whether to cache a specific content item only after it has been requested. This approach often relies on empirical formulas and classical caching algorithms, such as LRU (Least Recently Used) and LFU (Least Frequently Used), as well as their variants (Famaey et al., 2013; Shafiq et al., 2014). While these algorithms are simple and efficient, they frequently encounter challenges in selecting optimal parameter values, which can limit their performance in dynamic and diverse edge environments. (2) Proactive caching involves predictive methods to determine what content should be cached before any user requests are made. This approach leverages content popularity predictions and user behaviour profiles to make caching decisions at edge networks in advance. Advanced machine learning models, such as LSTM (Long Short-Term Memory) networks (Feng et al., 2019), are often employed to forecast content demand based on historical request patterns in proactive caching. These learning-driven methods generally offer superior caching performance compared to classical algorithms by automatically adjusting model parameters. However, they may require extensive computational capacity and high-quality training data, both of which may be often limited at the edge.
@@
Privacy concerns of edge caching. Despite the enormous benefits brought by caching content at the edge network, there has been some controversy over privacy violations brought by such caching. The concerns can be illustrated from two aspects. (1) The first privacy threat comes from external attackers, such as malicious user devices (Sivaraman and Sikdar, 2021; Liang and Liu, 2019; Acs et al., 2019; Qian et al., 2020; Cui et al., 2020a; Tong et al., 2023). The multi-access open property of the edge network provides an ideal entrance or interface for external attackers to obtain the cached content from the edge cache to extract sensitive information of end users (Xu et al., 2020). Adversaries can obtain user-sensitive information by launching cache side-channel attacks (Sivaraman and Sikdar, 2021; Liang and Liu, 2019; Acs et al., 2019) and cache tampering attacks (Qian et al., 2020; Cui et al., 2020a; Tong et al., 2023). However, it is non-trivial to embed advanced privacy protection mechanisms into edge networks due to the limited computing capacity, energy power, and storage space of edge devices. (2) Second, user privacy can be infringed by curious edge caching providers by analyzing traces and management records. Due to limited caching space relative to the rapidly growing user population and the scale of content (Zhou et al., 2019), edge network providers have a strong motivation to spy on user privacy in order to improve their resource utilization. In other words, if content popularity can be accurately predicted, the right content can be cached by edge devices just before the surge of requests towards content (Zhang et al., 2022d). Hence, edge network providers are curious about users’ personal interests and confidential information to infer their request behaviours, which can be extracted from users’ historical request traces (e.g., request patterns (Zhang et al., 2022d; Cui et al., 2020c, a), identifiable information (Zhang et al., 2022c; Zhu et al., 2021; Araldo et al., 2018; Cui et al., 2020c, a)). Edge network providers can implement monitoring attacks (Xue et al., 2018; Zhang et al., 2022c) and inference attacks (Qiao et al., 2022; Liu et al., 2022) in their systems to compromise users’ privacy based on collected request information from users. Therefore, an in-depth understanding of privacy risks in privacy-preserving edge caching (PPEC) is crucial for the design of feasible solutions to achieve privacy-preserving content cache at the edge network.
Our contributions. Recently, significant progress has been made in enhancing privacy protection for content caching at edge networks. However, these works fail to provide a comprehensive discussion of the privacy issues in edge caching systems. For instance, Ren et al. (Ren et al., 2019) primarily discuss the state-of-the-art researches on caching and privacy, respectively, in emerging edge computing paradigms. Besides, most surveys only discuss privacy issues for particular scenarios, such as IoT (Fotia et al., 2023; Sarwar et al., 2021), edge intelligence (Zhang et al., 2021), the metaverse (Otoum et al., 2024), and federated learning (FL) (Bellavista et al., 2021), while overlooking the distinct aspects of edge caching. The surveys in (Ni et al., 2021; Xiao et al., 2018) addressed privacy-preserving solutions and countermeasures for edge caching without covering all relevant issues in a thorough manner. In particular, there has been a lack of comprehensive discussion of protection methods targeting different types of private information in PPEC. Given these limitations and the absence of comprehensive literature reviews, this article aims to thoroughly examine and categorize current works on privacy issues in edge caching scenarios. The main contributions of this article are summarized as follows: (1) We make in-depth discussions on sensitive information in edge caching and propose a taxonomy from a private information perspective to classify existing works. To the best of our knowledge, this is the first such comprehensive exposition. (2) We conduct a thorough review of recent high-quality research, diving into the background of privacy attacks and mitigation methods in the realm of edge caching. Our review encompasses the latest solutions proposed for enhancing privacy in edge caching, which have been published in leading conferences and journals in the fields of computing networks, architecture, and privacy, such as CCS, INFOCOM, ToN, JSAC, TPDS, TIFS, and TDSC, as well as other top venues. Based on different kinds of privacy information and attacks towards each kind of privacy information, we respectively review countermeasures to defend against attacks for protecting each kind of infringed privacy. (3) Based on open problems outlined in existing works, we envision privacy-related open challenges in PPEC to provide insights for inspiring future research.